\begin{thebibliography}{14}
%
%\bibitem {A} T. Aoki, \textit{Calcul exponentiel des op\'erateurs
%microdifferentiels d'ordre infini.} I, Ann. Inst. Fourier (Grenoble)
%\textbf{33} (1983), 227--250.
%
%\bibitem {B} R. Brown, \textit{On a conjecture of Dirichlet},
%Amer. Math. Soc., Providence, RI, 1993.
%
%\bibitem {D} R. A. DeVore, \textit{Approximation of functions},
%Proc. Sympos. Appl. Math., vol. 36,
%Amer. Math. Soc., Providence, RI, 1986, pp. 34--56.

\bibitem{Adin2001}
R.M. Adin, F. Brenti, Y. Roichman, \textit{Descent numbers and major indices for the hyperoctahedral group},
Adv. in Appl. Math., \textbf{27} (2001), 210--224, .

%\bibitem{Athanasiadis14}
%C.A. Athanasiadis, \textit{Edgewise Subdivisions, local $h$-polynomials, and excedances in the
%wreath product $Z_r\wr\msn$}, SIAM J. Discrete Math., \textbf{28} (2014), 1479--1492.

\bibitem{Athanasiadis17}
C.A. Athanasiadis, \textit{Gamma-positivity in combinatorics and geometry}, S\'em. Lothar. Combin., \textbf{77} (2018), Article B77i.

%\bibitem{Athanasiadis18}
%C.A. Athanasiadis, \textit{Binomial Eulerian polynomials for colored permutations}, Arxiv:1812.00434.

\bibitem{Aval13}
J.-C. Aval, A. Boussicault, P. Nadeau, \textit{Tree-like tableaux}, Electron. J. Combin., \textbf{20}(4) (2013) Paper 34, 24 pp.

\bibitem{Bagno04}
E. Bagno, D. Garber, \textit{On the excedance number of colored permutation groups}, S\'em. Lothar. Combin., \textbf{53} (2004/2006), Article B53f.


\bibitem{Barbero14}
G.J.F. Barbero, J. Salas, E.J.S. Villase\~{n}or, \textit{Bivariate generating functions for a class of linear recurrences: general structure}, J. Combin. Theory Ser. A, \textbf{125} (2014), 146--165.
%\bibitem{Bagno07}
%E. Bagno and R. Biagioli, \textit{Colored descent representations for complex reflection groups}, Israel J. Math., \textbf{160} (2007), 317--347.

%\bibitem{Beck2010}
%M. Beck, A. Stapledon, \textit{On the log-concavity of Hilbert series of Veronese subrings and Ehrhart series}, Math. Z., \textbf{264} (2010), 195--207.
%
%\bibitem{Beck2013}
%M. Beck and B. Braun, \textit{Euler-Mahonian statistics via polyhedral geometry}, Adv. Math., \textbf{244} (2013), 925--954
%
%\bibitem{Beck2015}
%M. Beck and S. Robins, \textit{Computing the continuous discretely}, second ed., Springer Science Business Media, LLC, 2015.

\bibitem{Beck2019}
M. Beck, K. Jochemko and E. McCullough, \textit{$ h^\ast$-polynomials of zonotopes}, Trans. Amer. Math. Soc., \textbf{371} (2019), 2021--2042.

\bibitem{Bona08}
M. B\'ona, \textit{Real zeros and normal distribution for statistics on Stirling permutations defined by
Gessel and Stanley}, SIAM J. Discrete Math., \textbf{23} 2008/09, 401--406.

\bibitem{Borowiec16}
A. Borowiec, W. Mlotkowski, \textit{New Eulerian numbers of type $D$}, Electron. J. Combin., \textbf{23} (2016), \#P1.38.
%\bibitem{Branden04}
%P. Br\"and\'{e}n, \textit{Counterexamples to the Neggers-Stanley conjecture}, Electron. Res. Announc. Amer. Math. Soc., \textbf{10} (2004), 155--158.


%\bibitem{Branden06}
%P. Br\"and\'{e}n, \textit{On linear transformations preserving the P\'olya frequency property},
%Trans. Amer. Math. Soc., \textbf{358} (2006), 3697--3716.



%\bibitem{Branden08}
%P. Br\"and\'{e}n, \textit{Actions on permutations and unimodality of descent polynomials}, European J. Combin., \textbf{29} (2008), 514--531.

\bibitem{Branden11}
P. Br\"and\'{e}n, \textit{Iterated sequences and the geometry of zeros}, J. Reine Angew. Math., \textbf{658} (2011), 115--131.

\bibitem{Branden18}
P. Br\"and\'{e}n and L. Solus, \textit{Symmetric decompositions and real-rootedness}, Int Math. Res Notices, rnz059 (2019), https://doi.org/10.1093/imrn/rnz059.

%
\bibitem{Brenti90}
F. Brenti, \textit{Unimodal polynomials arising from symmetric functions}, Proc. Amer. Math. Soc., \textbf{108} (1990), 1133--1141.

\bibitem{Brenti94}
F. Brenti, \textit{$q$-Eulerian polynomials arising from Coxeter groups}, European J. Combin.,
\textbf{15} (1994), 417--441.

\bibitem{Brenti00}
F. Brenti, \textit{A class of $q$-symmetric functions arising from plethysm},
J. Combin. Theory Ser. A, \textbf{91} (2000), 137--170.

%\bibitem{Carlitz65}
%L. Carlitz, The coefficients in an asymptotic expansion, Proc. Amer. Math. Soc., 16 (1965), 248--252.

%\bibitem{Carlitz78}
%L. Carlitz and V.E. Hoggatt, \textit{Generalized Eulerian numbers and polynomials}, Fibonacci Quart., \textbf{16}(2) (1978), 138--146.

%\bibitem{Chen93}
%W.Y.C. Chen, \textit{Context-free grammars, differential operators and formal
%power series}, Theoret. Comput. Sci., \textbf{117} (1993), 113--129.
%
\bibitem{Chen09}
W.Y.C. Chen, R.L. Tang and A.F.Y. Zhao, \textit{Derangement polynomials and excedances of type $B$},
Electron. J. Combin., \textbf{16} (2) (2009), Research Paper 15.

%\bibitem{Chen16}
%W.Y.C. Chen, A.J.X. Guo, P.L. Guo, H.H.Y. Huang and T.Y.H. Liu, \textit{$\rm{s}$-inversion sequences and $P$-partitions of type $B$},
%SIAM J. Discrete Math., \textbf{30} (2016), 1632--1643.


%\bibitem{Chen17}
%W.Y.C. Chen, A.M. Fu, \textit{Context-free grammars for permutations and increasing trees}, Adv. in Appl. Math., \textbf{82} (2017), 58--82.

%\bibitem{Chen19}
%W.Y.C. Chen, H.R.L.Yang, \textit{A context-free grammar for the Ramanujan-Shor polynomials}, Adv. in Appl. Math., https://doi.org/10.1016/j.aam.2019.04.005.

\bibitem{Chow03}
C.-O. Chow, \textit{On the Eulerian polynomials of type $D$}, European J. Combin., \textbf{24} (2003), 391--408.


\bibitem{Chow08}
C.-O. Chow, \textit{On certain combinatorial expansions of the Eulerian polynomials}, Adv. in Appl. Math., \textbf{41} (2008), 133--157.

%\bibitem{Chow09}
%C.-O. Chow, \textit{On derangement polynomials of type $B$}, II, J. Combin. Theory Ser. A, \textbf{116} (2009),
%816--830.


\bibitem{Chow14}
C.-O. Chow, S.-M. Ma, \textit{Counting signed permutations by their alternating runs}, Discrete Math., \textbf{323} (2014), 49--57.
%
%\bibitem{Chow10}
%C.-O. Chow and T. Mansour, \textit{Counting derangements, involutions and unimodal elements in the
%wreath product $C_r\wr \msn$}, Israel J. Math., \textbf{179} (2010), 425--448.

%\bibitem{Diaconis14}
%P. Diaconis, S.N. Evans, R. Graham, \textit{Unseparated pairs and fixed points in random permutations}, Adv. in Appl. Math., \textbf{61} (2014), 102--124.

\bibitem{Taylor03}
M.A. Eisenstein-Taylor, Polytopes, \textit{permutation shapes and bin packing}, Adv. Appl. Math., \textbf{30} (2003), 96--109.

%\bibitem{Duh18}
%G.-H. Duh, Y.-C. Roger Lin, S.-M. Ma, Y.-N. Yeh, \textit{Some statistics on Stirling permutations and Stirling derangements}, Discrete Math., \textbf{341} (2018), 2478--2484.


%\bibitem{Dumont96}
%D. Dumont, \textit{Grammaires de William Chen et d\'erivations dans les arbres et
%arborescences}, S\'em. Lothar. Combin., \textbf{37}, Art. B37a (1996), 1--21.

%\bibitem{Ehrhart62}
%E. Ehrhart, \textit{Sur les poly\`{e}dres rationnels homoth\'{e}tiques \`{a} $n$ dimensions}, C. R. Acad. Sci. Paris, \textbf{254} (1962), 616--618.
\bibitem{Foata11}
D. Foata, G.-N Han, \textit{The decrease value theorem with an application
to permutation statistics}, Adv. in Appl. Math., \textbf{46} (2011), 296--311.


%\bibitem{Foata70}
%D. Foata, M. P. Sch\"utzenberger, \textit{Th\'eorie g\'eometrique des polyn\^omes eul\'eriens}, Lecture Notes in Math., vol. \textbf{138}, Springer, Berlin, 1970.
%
%\bibitem{Foata73}
%D. Foata, V. Strehl, \textit{Euler numbers and variations of permutations}, in Colloquio Internazionale
%sulle Teorie Combinatorie (Roma, 1973), Tomo I, Atti dei Convegni Lincei, No. \textbf{17}, Accad. Naz.
%Lincei, Rome, 1976, pp. 119--131.

%\bibitem{Foata11}
%D. Foata and G.-N. Han, \textit{The decrease value theorem with an application to permutation statistics},
%Adv. Appl. Math., \textbf{46} (2011), 296--311.
%
%\bibitem{Gasharov98}
%V. Gasharov, \textit{On the Neggers-Stanley conjecture and the Eulerian polynomials}, J. Combin. Theory Ser. A,  \textbf{82} (2) (1998), 134--146.
%
%\bibitem{Haglund99}
%J. Haglund, K. Ono, D.G. Wagner, \textit{ Theorems and conjectures involving rook polynomials with only real zeros}, In: Ahlgren S.D., Andrews G.E., Ono K. (eds) Topics in Number Theory. Mathematics and Its Applications, vol \textbf{467}. (1999), Springer, Boston, MA.


%\bibitem{Fu18}
%A.M. Fu, \textit{A context-free grammar for peaks and double descents of permutations}, Adv. in Appl. Math., \textbf{100} (2018), 179--196.

\bibitem{Gal05}
S.R. Gal, \textit{Real root conjecture fails for five and higher-dimensional spheres}, Discrete Comput. Geom., \textbf{34} (2005), 269--284.

\bibitem{Gessel78}
I. Gessel and R.P. Stanley, \textit{Stirling polynomials}, J. Combin. Theory Ser.
A, \textbf{24} (1978), 25--33.

\bibitem{Haglund12}
J. Haglund, M. Visontai, \textit{Stable multivariate Eulerian polynomials and generalized Stirling permutations}, European J. Combin., \textbf{33} (2012), 477--487.

\bibitem{Hwang20}
H.-K. Hwang, H.-H. Chern, G.-H. Duh, \textit{An asymptotic distribution theory for Eulerian
recurrences with applications}, Adv. in Appl. Math., \textbf{112} (2020), 101960.

\bibitem{Hyatt16}
M. Hyatt, \textit{Recurrences for Eulerian polynomials of type $B$ and type $D$}, Ann. Comb., \textbf{20} (2016), 869--881.


%\bibitem{Gessel18}
%I. Gessel, Y. Zhuang, \textit{Shuffle-compatible permutation statistics}, Adv. Math, \textbf{332} (2018), 85--141.
%
%\bibitem{Janson11}
%S. Janson, M. Kuba and A. Panholzer, \textit{Generalized Stirling permutations, families of increasing trees and urn models},
%J. Combin. Theory Ser. A, \textbf{118} (2011), 94--114.



%\bibitem{Lin1501}
%Z. Lin, \textit{On the descent polynomial of signed multipermutations}, Proc. Amer. Math. Soc., \textbf{143} (9) (2015), 3671--3685.

%\bibitem{Lin15}
%Z. Lin, J. Zeng, \textit{The $\gamma$-positivity of basic Eulerian polynomials via group actions}, J. Combin. Theory Ser. A, \textbf{135} (2015), 112--129.

%\bibitem{Linusson12}
%S. Linusson, J. Shareshian and M.L. Wachs, \textit{Rees products and lexicographic shellability},  J. Comb., \textbf{3} (2012), 243--276.

%
%\bibitem{Petersen07}
%T.K. Petersen, Enriched P-partitions and peak algebras, Adv. Math. 209 (2007), 561--610.
\bibitem{Liu07}
L.L. Liu, Y. Wang, \textit{A unified approach to polynomial sequences with only real zeros}, Adv. in Appl. Math., \textbf{38} (2007), 542--560.

\bibitem{Ma121}
S.-M. Ma, \textit{Derivative polynomials and enumeration of permutations by number of interior and left peaks}, Discrete Math., \textbf{312} (2012), 405--412.

\bibitem{Ma132}
S.-M. Ma, \textit{Enumeration of permutations by number of alternating runs}, Discrete Math., \textbf{313} (2013), 1816--1822.
%

\bibitem{Ma15}
S.-M. Ma, T. Mansour, \textit{The $1/k$-Eulerian polynomials and $k$-Stirling permutations}, Discrete Math., \textbf{338} (2015), 1468--1472.

\bibitem{Ma17}
S.-M. Ma, Y.-N. Yeh, \textit{Eulerian polynomials, Stirling permutations
of the second kind and perfect matchings}, Electron. J. Combin.,
\textbf{24}(4) (2017), \#P4.27.


%\bibitem{Ma1801}
%S.-M. Ma, J. Ma, Y.-N. Yeh, B.-X. Zhu, \textit{Context-free grammars for several polynomials
%associated with Eulerian polynomials}, Electron. J. Combin.,
%\textbf{25}(1) (2018), \#P1.31.

\bibitem{Ma19}
S.-M. Ma, J. Ma, Y.-N. Yeh, \textit{$\gamma$-positivity and partial $\gamma$-positivity of descent-type polynomials}, J. Combin. Theory Ser. A, \textbf{167} (2019), 257--293.

\bibitem{Ma2001}
S.-M. Ma, J. Ma, Y.-N. Yeh, \textit{David-Barton type identities and the alternating run polynomials}, Adv. in Appl. Math., \textbf{114} (2020), 101978.


\bibitem{Ma2008}
S.-M. Ma, Y. Wang, \textit{$q$-Eulerian polynomials and polynomials with only real zeros}, Electron. J. Combin., \textbf{15} (2008), \#R17.

%
%\bibitem{MacMahon20}
%P.A. MacMahon, \textit{Combinatory Analysis}, Vol. I, II (bound in one volume), Dover Phoenix
%Editions, Dover Publications Inc., Mineola, NY, 2004. Reprint of An Introduction to Combinatory
%Analysis (1920) and Combinatory Analysis. Vol. I, II (1915, 1916).

\bibitem{Mongelli13}
P. Mongelli, \textit{Excedances in classical and affine Weyl groups}, J. Combin. Theory Ser. A, \textbf{120} (2013), 1216--1234.

\bibitem{Petersen06}
T.K. Petersen, \textit{Descents, peaks, and $P$-partitions}, Ph.D. Dissertation, Brandeis University, 2006.

\bibitem{Petersen07}
T.K. Petersen, \textit{Enriched $P$-partitions and peak algebras}, Adv. Math., \textbf{209}(2) (2007), 561--610.

%\bibitem{Petersen15}
%T.K. Petersen, \textit{Eulerian Numbers}. Birkh\"auser/Springer, New York, 2015.

%\bibitem{Roselle68}
%D.P. Roselle, \textit{Permutations by number of rises and successions}, Proc. Amer. Math. Soc., \textbf{19} (1968), 8--16.


%\bibitem{Savage12}
%C.D. Savage and G. Viswanathan, \textit{The $1/k$-Eulerian polynomials}, Electron J. Combin.,
%\textbf{19} (2012), \#P9.

%\bibitem{Savage1202}
%C.D. Savage and Michael J. Schuster, \textit{Ehrhart series of lecture hall polytopes and Eulerian
%polynomials for inversion sequences}, J. Combin. Theory Ser. A, \textbf{119}(4) (2012), 850--870, .

\bibitem{Savage15}
C.D. Savage and M. Visontai, \textit{The $s$-Eulerian polynomials have only real roots}, Trans. Amer. Math.
Soc., \textbf{367} (2015), 763--788.

\bibitem{Schepers13}
J. Schepers and L.V. Langenhoven, \textit{Unimodality questions for integrally closed lattice polytopes}, Ann.
Comb., \textbf{17}(3) (2013), 571--589.

%\bibitem{Shareshian09}
%J. Shareshian and M.L. Wachs, \textit{Poset homology of Rees products and $q$-Eulerian polynomials}, Electron. J. Combin., \textbf{16}(2) (2009), R20.

%\bibitem{Zeng12}
%H. Shin, J. Zeng, \textit{The symmetric and unimodal expansion of Eulerian polynomials via continued fractions}, European J. Combin., \textbf{33} (2012), 111--127.

\bibitem{Zeng16}
H. Shin and J. Zeng, \textit{Symmetric unimodal expansions of excedances in colored permutations}, European
J. Combin., \textbf{52} (2016), 174--196.

\bibitem{Sloane}
N.J.A. Sloane, The On-Line Encyclopedia of Integer Sequences,
published electronically at http://oeis.org, 2010.

%\bibitem{Simion84}
%R. Simion, \textit{A multi-indexed Sturm sequence of polynomials and unimodality of certain
%combinatorial sequences}, J. Combin. Theory Ser. A, \textbf{36} (1984), no. 1, 15--22,

\bibitem{Solus19}
L. Solus, \textit{Simplices for numeral systems}, Trans. Amer. Math. Soc., \textbf{371} (2019), 2089--2107.

\bibitem{Stanley08}
R.P. Stanley, \textit{Longest alternating subsequences of permutations}, Michigan Math. J., \textbf{57} (2008), 675--687.
%
%\bibitem{Stanley92}
%R.P. Stanley, \textit{Subdivisions and local $h$-vectors}, J. Amer. Math. Soc., \textbf{5} (1992), 805--851.

%\bibitem{Stapledon09}
%A. Stapledon, \textit{Inequalities and Ehrhart $\delta$-vectors}, Trans. Amer. Math. Soc., \textbf{361} (2009), 5615--5626.

%\bibitem{Steingrimsson}
%E. Steingr\'imsson, \textit{Permutation statistics of indexed permutations}, European J. Combin.,
%\textbf{15} (1994), 187--205.

%\bibitem{Stembridge92}
%J.R. Stembridge, \textit{J.R. Stembridge, Eulerian numbers, tableaux, and the Betti numbers of a toric variety}, Discrete
%Math., \textbf{99} (1992), 307--320.

\bibitem{Stembridge94}
J.R. Stembridge, \textit{Some permutation representations of Weyl groups associated with the cohomology of toric varieties},
Adv. Math., \textbf{106} (1994), 244--301.

\bibitem{Wang2005}
Y. Wang, Y.-N. Yeh, \textit{Proof of a conjecture on unimodality}, Eur. J. Combin., \textbf{26} 2005), 617--627.

%\bibitem{Visontai13}
%M. Visontai, N. Williams, Stable multivariate W-Eulerian polynomials,
%J. Combin. Theory Ser. A 120 (2013), 1929--1945.


%\bibitem{Stembridge}
%J.R. Stembridge, \textit{Some permutation representations of Weyl groups assoiated with the cohomology of tori varieties}, Adv. Math., \textbf{106}(2) (1994), 244--301.
%%\bibitem{Stembridge07}
%
%J.R. Stembridge, \textit{Counterexamples to the poset conjectures of Neggers, Stanley, and Stembridge}, Trans. Amer. Math. Soc., \textbf{359} (3) (2007), 1115--1128.

\bibitem{Yang}
A.L.B. Yang and P.B. Zhang, \textit{Brenti's open problem on the real-rootedness of $q$-Eulerian polynomials of type $D$}, SIAM J. Discrete Math., \textbf{31}(2), 918--926.
%

\bibitem{Zhao11}
A.F.Y. Zhao, \textit{The combinatorics on permutations and derangements of type $B$}, Ph.D. dissertation, Nankai University, 2011.

\bibitem{Zhu}
B.-X.Zhu, \textit{A generalized Eulerian triangle from staircase tableaux and tree-like tableaux}, J. Combin. Theory Ser. A, \textbf{172} (2020), 105206.


\bibitem{Zhuang17}
Y. Zhuang, \textit{Eulerian polynomials and descent statistics}, Adv. in Appl. Math., \textbf{90} (2017), 86--144.
\end{thebibliography}

\documentclass[11pt,reqno]{amsart}

\usepackage{graphicx}

\usepackage{xspace}
\usepackage{amsmath,amsthm,amsfonts,amssymb,latexsym,mathrsfs,color}
\usepackage{hyperref}

\newtheorem{theorem}{Theorem}
\newtheorem{corollary}[theorem]{Corollary}
\newtheorem{proposition}[theorem]{Proposition}

\newtheorem{definition}[theorem]{Definition}

\newcommand{\negg}{{\rm neg\,}}

\newcommand{\altrun}{{\rm altrun\,}}

\newcommand{\udrun}{{\rm udrun\,}}
\newcommand{\val}{{\rm val\,}}
\newcommand{\fap}{{\rm fap\,}}

\newcommand{\fdes}{{\rm fdes\,}}

\newcommand{\ap}{{\rm ap\,}}
\newcommand{\lap}{{\rm lap\,}}
\newcommand{\lpk}{{\rm lpk\,}}

\newcommand{\ipk}{{\rm ipk\,}}
\newcommand{\pk}{{\rm pk\,}}

\newcommand{\des}{{\rm des\,}}

\newcommand{\exc}{{\rm exc\,}}
\newcommand{\fexc}{{\rm fexc\,}}

\newcommand{\mdn}{\mathcal{S}^D_n}

\newcommand{\msn}{\mathcal{S}_n}

\newcommand{\lrf}[1]{\lfloor #1\rfloor}

\newcommand{\mbn}{{\mathcal S}^B_n}

\newcommand{\mqn}{\mathcal{Q}_n}

\DeclareMathOperator{\LHP}{LHP}
\DeclareMathOperator{\R}{\mathbb{R}}
\DeclareMathOperator{\re}{Re}

\newcommand{\eulerian}[2]{\genfrac{|}{|}{0pt}{}{#1}{#2}}

\title{Eulerian pairs and Eulerian recurrence systems}
\author[S.-M.~Ma]{Shi-Mei Ma}
\address{School of Mathematics and Statistics,
        Northeastern University at Qinhuangdao,
         Hebei 066000, P.R. China}
\email{shimeimapapers@163.com (S.-M. Ma)}
\author[J.~Ma]{Jun Ma}
\address{Department of mathematics, Shanghai Jiao Tong University, Shanghai, P.R. China}
\email{majun904@sjtu.edu.cn(J.~Ma)}
\author{Jean Yeh}
\address{Department of Mathematics, National Kaohsiung Normal University, Kaohsiung 82444, Taiwan}
\email{chunchenyeh@nknu.edu.tw}
\author[Y.-N. Yeh]{Yeong-Nan Yeh}
\address{Institute of Mathematics,
        Academia Sinica, Taipei, Taiwan}
\email{mayeh@math.sinica.edu.tw (Y.-N. Yeh)}
\thanks{The first and third authors are both corresponding authors}
\subjclass[2010]{Primary 05A05; Secondary 26C05}
\begin{document}

\maketitle
\begin{abstract}
In this paper, we characterize a duality relation between Eulerian recurrences and Eulerian recurrence systems,
which generalizes and unifies Hermite-Biehler decompositions of several enumerative polynomials, including flag descent polynomials for hyperoctahedral group, flag ascent-plateau polynomials for Stirling permutations, up-down run polynomials for symmetric group and alternating run polynomials for hyperoctahedral group.
As applications, we derive some properties of associated enumerative polynomials.
In particular, we find that both the ascent-plateau polynomials and left ascent-plateau polynomials for Stirling permutations are alternatingly increasing,
and so they are unimodal with modes in the middle.

\bigskip

\noindent{\sl Keywords}: Eulerian polynomials, Eulerian pairs, Eulerian recurrence systems
\end{abstract}
\date{\today}
\section{Introduction}
Given $m,n\in \mathbb{N}$. When $m\leq n$, let $[m,n]=\{m,m+1,\ldots,n\}$. As usual, let $[n]=\{1,2,\ldots,n\}$.
The cardinality of a set $A$ will be
denoted by $\#A$. Let $\msn$ be the set of all permutations of $[n]$ and
let $\pi=\pi(1)\pi(2)\cdots\pi(n)\in\msn$.
Let $\mbn$ be the hyperoctahedral group of rank $n$. Elements of $\mbn$ are signed permutations $\pi$ of the set $\pm[n]$ such that $\pi(-i)=-\pi(i)$ for all $i$, where $\pm[n]=\{\pm1,\pm2,\ldots,\pm n\}$.
Let $\mdn$ denote the group of even signed permutations, which is a Coxeter group of type $D$ of rank $n$.
The group $\mdn$ is the subgroup of $\mbn$ consisting of signed permutations with an even number of negative entries among $\pi(1),\pi(2)\ldots,\pi(n)$.
Define
\begin{equation*}
\begin{split}
\des_A(\pi)&:=\#\{i\in [n-1]:~\pi(i)>\pi({i+1})\},\\
\des_B(\pi)&:=\#\{i\in [0,n-1]:~\pi(i)>\pi({i+1}),~\pi(0)=0\},\\
\des_D(\pi)&:=\#\{i\in[0,n-1]:~\pi(i)>\pi({i+1}),~\pi(0)=-\pi(2)\}.
\end{split}
\end{equation*}
The types $A,B$ and $D$ Eulerian polynomials are
respectively defined by
\begin{equation*}
\begin{split}
A_n(x)&=\sum_{\pi\in\msn}x^{\des_A(\pi)},~
B_n(x)=\sum_{\pi\in\mbn}x^{\des_B(\pi)},~
D_n(x)=\sum_{\pi\in \mdn}x^{\des_D(\pi)}.
\end{split}
\end{equation*}
There is a close connection among the three types of Eulerian polynomials~\cite[Lemma 9.1]{Stembridge94}:
\begin{equation}\label{Dnx-BnxAnx}
D_n(x)=B_n(x)-n2^{n-1}xA_{n-1}(x) \quad {\text for}\quad n\geq 2.
\end{equation}
Subsequently, Brenti~\cite[Theorem~4.7]{Brenti94} obtained a $q$-analogue of~\eqref{Dnx-BnxAnx}. There are several recurrences for $D_n(x)$ (see~\cite{Chow03,Hyatt16}).
The polynomials $A_n(x),B_n(x)$ and $D_n(x)$ have several common properties, including unimodality, real-rootedness and $\gamma$-positivity
(see~\cite{Athanasiadis17,Hyatt16,Liu07,Petersen07,Savage15,Yang}).
This paper is motivated by the recent work of Hyatt~\cite{Hyatt16} and Hwang, Chern and Duh~\cite{Hwang20}.

Let $\mbn=B_n^+\cup B_n^-$ and $\mdn=D_n^+\cup D_n^-$, where
$$B_n^+=\{\pi\in\mbn: \pi(n)>0\},~B_n^-=\{\pi\in\mbn:~\pi(n)<0\},$$
$$D_n^+=\{\pi\in\mdn: \pi(n)>0\},~D_n^-=\{\pi\in\mdn:~\pi(n)<0\}.$$
Define $$P_n(x)=\sum_{k=0}^{n-1}\binom{n}{k}B_k(x)(x-1)^{n-k-1},~Q_n(x)=\sum_{k=0}^{n-1}\binom{n}{k}D_k(x)(x-1)^{n-k-1}.$$
Given a polynomial $f(x)$ of degree $n$. Let $\widetilde{f}(x)=x^nf(1/x)$.
In~\cite{Hyatt16}, Hyatt found that
\begin{equation}\label{Bnx}
B_n(x)=P_n(x)+x^nP_n(1/x)=P_n(x)+x\widetilde{P}_n(x)~\text{for $n\geq 1$},
\end{equation}
\begin{equation*}\label{Dnx}
D_n(x)=Q_n(x)+x^nQ_n(1/x)=Q_n(x)+x\widetilde{Q}_n(x)~\text{for $n\geq 2$},
\end{equation*}
where
\begin{equation*}
P_n(x)=\sum_{\pi\in B_n^+}x^{\des_B(\pi)},~x^nP_n(1/x)=\sum_{\pi\in B_n^-}x^{\des_B(\pi)},
\end{equation*}
\begin{equation*}
Q_n(x)=\sum_{\pi\in D_n^+}x^{\des_D(\pi)},~x^nQ_n(1/x)=\sum_{\pi\in D_n^-}x^{\des_D(\pi)}.
\end{equation*}
Motivated by~\eqref{Bnx}, in the following we shall present a more natural decomposition of $B_n(x)$.

Following~\cite{Adin2001}, the {\it flag descent number} of $\pi\in\mbn$
is defined by
\begin{equation*}
\fdes(\pi):=\begin{cases}
2\des_A(\pi)+1,& \text{if $\pi(1)<0$};\\
2\des_A(\pi), & \text{otherwise}.
\end{cases}
\end{equation*}
Clearly, $\fdes(\pi)=\des_A(\pi)+\des_B(\pi)$. The {\it flag descent polynomial} is defined by
$$C_n(x)=\sum_{\pi\in \mbn}x^{\fdes(\pi)}.$$
It follows from~\cite[Theorem~4.4]{Adin2001} that
\begin{equation}\label{fdesAnx}
C_n(x)=(1+x)^nA_n(x),
\end{equation}
and so $C_n(x)$ is symmetric and unimodal. Let $\negg(\pi):=\#\{i\in [n]:~\pi(i)<0\}$.
Consider the $q$-flag descent polynomials
$$C_n(x,q)=\sum_{\pi\in \mbn}x^{\fdes(\pi)}q^{\negg(\pi)}.$$
Set $\mbn=C_n^+\cup C_n^-$, where $C_n^+=\{\pi\in\mbn:~\pi(1)>0\}$ and $C_n^-=\{\pi\in \mbn:~\pi(1)<0\}$.
Define
$$C_n^E(x,q)=\sum_{\pi\in C_n^+}x^{\des_A(\pi)}q^{\negg(\pi)},~C_n^O(x,q)=\sum_{\pi\in C_n^-}x^{\des_A(\pi)}q^{\negg(\pi)}.$$
Then we have
\begin{equation}\label{SnxqEO}
C_n(x,q)=C_n^E(x^2,q)+xC_n^O(x^2,q).
\end{equation}

Consider the {\it type $B$ $q$-Eulerian polynomials}
$$B_n(x,q)=\sum_{\pi\in \mbn}x^{\des_B(\pi)}q^{\negg(\pi)}.$$
The polynomials $B_n(x,q)$ satisfy the recurrence
$$B_{n+1}(x,q)=((1+q)nx+qx+1)B_n(x,q)+(1+q)x(1-x)B_n'(x,q),$$
with $B_0(x,q)=1$ (see~\cite[Theorem~3.4]{Brenti94}).
Note that $\des_B(\pi)=\des_A(\pi)$ for $\pi\in C_n^+$ and $\des_B(\pi)=\des_A(\pi)+1$ for $\pi\in C_n^-$.
Hence
\begin{equation*}
B_n(x,q)=C_n^E(x,q)+xC_n^O(x,q).
\end{equation*}
Set $C_n^E(x,1)=C_n^E(x)$ and $C_n^O(x,1)=C_n^O(x)$. By~\eqref{fdesAnx}, the polynomial $C_n(x)$ is symmetric. Thus $C_n^E(x)=x^{n-1}C_n^O(1/x)$ for $n\geq 1$.
Note that $B_n(x,1)=B_n(x)$ and $C_n(x,1)=C_n(x)$. We can now conclude the following result from the discussion above.
\begin{proposition}\label{SnxBnx-prop}
For $n\geq 1$, we have
\begin{equation}\label{SnxBnx-decom}
C_n(x)=C_n^E(x^2)+xC_n^O(x^2),~B_n(x)=C_n^E(x)+xC_n^O(x).
\end{equation}
\end{proposition}

Comparing~\eqref{Bnx} with~\eqref{SnxBnx-decom}, we get the following result, and we give a proof of it for completeness.
\begin{proposition}
We have $C_n^E(x)=P_n(x)$ and $C_n^O(x)=\widetilde{P}_n(x)$.
\end{proposition}
\begin{proof}
Define
\begin{align*}
{^+B_n^+}&=\{\pi\in B_n: \pi(1)>0,~ \pi(n)>0\},\\
{^+B_n^-}&=\{\pi\in B_n: \pi(1)>0,~ \pi(n)<0\},\\
{^-B_n^+}&=\{\pi\in B_n: \pi(1)<0,~ \pi(n)>0\},\\
{^-B_n^-}&=\{\pi\in B_n: \pi(1)<0,~ \pi(n)<0\}.
\end{align*}
Note that $C_n^+={^+B_n^+}\cup {^+B_n^-}$ and $B_n^+={^+B_n^+}\cup {^-B_n^+}$.
A bijection $\Phi$ from $C_n^+$ to $B_n^+$ is given as follows:
\begin{itemize}
  \item [$(i)$] If $\pi\in {^+B_n^+}$, then let $\Phi(\pi)=\pi$;
  \item  [$(ii)$] For $\pi\in {^+B_n^-}$, let $k$ be the smallest index of $\pi$ such that $\pi(k)>0$ and $\pi(k+1)<0$.
  Then we define $\Phi(\pi)=\pi(k+1)\cdots \pi(n)\pi(1)\cdots\pi(k)$.
\end{itemize}
Note that $\des_B\left(\Phi(\pi)\right)=\des_B(\pi)$. Hence
\begin{equation*}
\sum_{\pi\in C_n^+}x^{\des_B(\pi)}=\sum_{\pi\in B_n^+}x^{\des_B(\pi)},
\end{equation*}
Along the same lines, it is easy to show that
$\sum_{\pi\in C_n^-}x^{\des_B(\pi)}=\sum_{\pi\in B_n^-}x^{\des_B(\pi)}$.
\end{proof}

Motivated by Proposition~\ref{SnxBnx-prop}, in the following we shall establish a duality relation between Eulerian recurrences and Eulerian recurrence systems.
It is well known that the Eulerian polynomials $A_n(x)$ and $B_n(x)$ satisfy the following recurrence relations:
\begin{equation*}
\begin{split}
A_{n}(x)&=(nx+1-x)A_{n-1}(x)+x(1-x)A_{n-1}'(x),\\
B_{n}(x)&=(2nx+1-x)B_{n-1}(x)+2x(1-x)B_{n-1}'(x),
\end{split}
\end{equation*}
with the initial conditions $A_0(x)=B_0(x)=1$ (see~\cite{Brenti94,Zhuang17}).
In recent years, there has been much work on the generalizations of Eulerian recurrences, see~\cite{Barbero14,Hwang20,Zhu} and references therein.
In~\cite{Barbero14}, Barbero, Salas and Villase\~{n}or systematically studied and classified the partial differential equations that are satisfied by
the generating function
$$f(x,y)=\sum_{n,k\geq 0}\eulerian{n}{k}x^k\frac{y^n}{n!},$$
where the numbers $\eulerian{n}{k}$ satisfy the recurrence relation
$$\eulerian{n}{k}=(\alpha n+\beta k+\gamma)\eulerian{n-1}{k-1}+(\alpha' n+\beta' k+\gamma')\eulerian{n-1}{k},$$
with $\eulerian{0}{0}=1$ and $\eulerian{n}{k}=0$ when $n<0$ or $k<0$.
Very recently, Hwang, Chern and Duh~\cite{Hwang20} considered the general Eulerian recurrence:
\begin{equation}\label{Eulerian02}
P_{n}(x)=(\alpha(x)n+\gamma(x))P_{n-1}(x)+\beta(x)(1-x)P_{n-1}'(x),
\end{equation}
with $P_0(x),\alpha(x),\beta(x)$ and $\gamma(v)$ are given functions of $x$ (they are often polynomials). They studied the limiting distribution of the coefficients of $P_n(x)$ for large $n$ when the coefficients are nonnegative. In particular, Hwang, Chern and Duh~\cite[Section~9.3]{Hwang20} discussed the limiting distribution of the coefficients polynomials that satisfy Eulerian recurrence systems.

\begin{definition}\label{def01}
Let $\{E_n(x)\}_{n\geq 0}$ and $\{O_n(x)\}_{n\geq 0}$ be two sequences of polynomials.
We say that the ordered pair of polynomials $(E_n(x),O_n(x))$ is a Eulerian pair if $\deg E_n(x)\geq \deg O_n(x)$ and
the polynomials $E_n(x)$ and $O_n(x)$ satisfy the Eulerian recurrence system:
\begin{equation}\label{EulerianPair}
\left\{
  \begin{array}{l}
    E_{n+1}(x)=p_n(x)E_n(x)+q_n(x)E_n'(x)+r_n(x)O_n(x), \\
    O_{n+1}(x)=u_n(x)O_n(x)+v_n(x)O_n'(x)+w_n(x)E_n(x),
  \end{array}
\right.
\end{equation}
where $E_0(x),O_0(x),p_n(x),q_n(x),r_n(x),u_n(x),v_n(x),w_n(x)$ are given polynomials of low degrees.
\end{definition}

Let $f(x)=\sum_{i=0}^nf_ix^i$. Throughout this paper, we always let
\begin{align*}
f^E(x)&=\sum_{k=0}^{\lrf{\frac{n}{2}}}f_{2k}x^{k},~f^O(x)=\sum_{k=0}^{\lrf{\frac{n-1}{2}}}f_{2k+1}x^{k};\\
f^e(x)&=\sum_{k=0}^{\lrf{\frac{n}{2}}}f_{2k}x^{2k},~f^o(x)=\sum_{k=0}^{\lrf{\frac{n-1}{2}}}f_{2k+1}x^{2k+1}.
\end{align*}
Hence $f(x)=f^E(x^2)+xf^O(x^2)=f^e(x)+f^o(x)$.
As an extension of~\eqref{Eulerian02}, we define a sequence of polynomials $F_n(x)$ by using the following general Eulerian recurrence:
\begin{equation}\label{Eulerian03}
F_{n+1}(x)=\alpha_n(x)F_{n}(x)+\beta_n(x)F_{n}'(x),
\end{equation}
where $F_0(x),\alpha_n(x)$ and $\beta_n(x)$ are given polynomials of low degrees.
We can now present the first main result of this paper.
\begin{theorem}\label{mathm01}
Let $\langle E_n(x),O_n(x)\rangle$ be a Eulerian pair that satisfies the Eulerian recurrence syetem~\eqref{EulerianPair},
and let $F_n(x)$ be the polynomial defined by the recurrence~\eqref{Eulerian03}.
Then the polynomial $F_n(x)$ has the expression $F_n(x)=E_n(x^2)+xO_n(x^2)$ if and only if the following conditions hold:
\begin{align*}
u_n(x)&=p_n(x)+\frac{1}{2x}q_n(x),~v_n(x)=q_n(x),~w_n(x)=\frac{1}{x}r_n(x),\\
\alpha_n(x)&=p_n(x^2)+\frac{1}{x}r_n(x^2),~\beta_n(x)=\frac{1}{2x}q_n(x^2),~\beta_n^e(x)=0.
\end{align*}
\end{theorem}
\begin{proof}
By using $F_n(x)=E_n(x^2)+xO_n(x^2)$, we obtain
\begin{equation*}
F_n'(x)=2xE_n'(x^2)+O_n(x^2)+2x^2O_n'(x^2).
\end{equation*}
Then it follows from~\eqref{Eulerian03} that
\begin{align*}
F_{n+1}(x)&=\alpha_n(x)\left(E_n(x^2)+xO_n(x^2)\right)+\\
&\beta_n(x)\left(2xE_n'(x^2)+O_n(x^2)+2x^2O_n'(x^2)\right).
\end{align*}
Comparing this with the expression
$F_{n+1}(x)=E_{n+1}(x^2)+xO_{n+1}(x^2)$, we obtain
\begin{align*}
E_{n+1}(x^2)&=\alpha_n^e(x)E_n(x^2)+x\alpha_n^o(x)O_n(x^2)+\\
&\beta_n^e(x)\left(O_n(x^2)+2x^2O_n'(x^2)\right)+2x\beta_n^o(x)E_n'(x^2),\\
O_{n+1}(x^2)&=\frac{1}{x}\alpha_n^o(x)E_n(x^2)+\alpha_n^e(x)O_n(x^2)+\\
&\frac{1}{x}\beta_n^o(x)\left(O_n(x^2)+2x^2O_n'(x^2)\right)+2\beta_n^e(x)E_n'(x^2).
\end{align*}
Since~\eqref{EulerianPair} holds, then $\beta_n^e(x)=0$.
Hence
\begin{equation}\label{EulerianPair02}
\left\{
  \begin{array}{l}
E_{n+1}(x^2)=\alpha_n^e(x)E_n(x^2)+2x\beta_n^o(x)E_n'(x^2)+x\alpha_n^o(x)O_n(x^2),\\
O_{n+1}(x^2)=\left(\alpha_n^e(x)+\frac{1}{x}\beta_n^o(x)\right)O_n(x^2)+2x\beta_n^o(x)O_n'(x^2)+\frac{1}{x}\alpha_n^o(x)E_n(x^2).
  \end{array}
\right.
\end{equation}
By comparing~\eqref{EulerianPair} with~\eqref{EulerianPair02}, we immediately get the following relations:
\begin{align*}
p_n(x^2)&=\alpha_n^e(x),~q_n(x^2)=2x\beta_n^o(x),~r_n(x^2)=x\alpha_n^o(x),\\
u_n(x^2)&=\alpha_n^e(x)+\frac{1}{x}\beta_n^o(x),~v_n(x^2)=2x\beta_n^o(x),~w_n(x^2)=\frac{1}{x}\alpha_n^o(x),
\end{align*}
which yield the desired result. This completes the proof.
\end{proof}

Let $F_n(x)$ be the polynomials defined by the recurrence~\eqref{Eulerian03}.
When $\beta_n^e(x)=0$, we can get the recurrence system~\ref{EulerianPair} by using
the following relations:
\begin{equation}\label{dualrelation01}
x\alpha_n(x)=xp_n(x^2)+r_n(x^2),~2x\beta_n(x)=q_n(x^2).
\end{equation}

A polynomial $p\in\R[x]$ is ({\it Hurwitz} or {\it asymptotically})
{\it stable} if every zero of $p$ is in the open left half plane
$\LHP=\{z\colon\re z<0\}$;
$p$ is {\it standard} if its leading coefficient is positive.
Suppose that $p,q\in\R[x]$ both have only real zeros,
that those of $p$ are $\xi_1\leqslant\cdots\leqslant\xi_n$,
and that those of $q$ are $\theta_1\leqslant\cdots\leqslant\theta_m$.
We say that $p$ {\it interlaces} $q$ if $\deg q=1+\deg p$ and the zeros of
$p$ and $q$ satisfy
$$
\theta_1\leqslant\xi_1\leqslant\theta_2\leqslant\cdots\leqslant\xi_n
\leqslant\theta_{n+1}.
$$
We also say that $p$ {\it alternates left of} $q$ if $\deg p=\deg q$
and the zeros of $p$ and $q$ satisfy
$$
\xi_1\leqslant\theta_1\leqslant\xi_2\leqslant\cdots\leqslant\xi_n
\leqslant\theta_n.
$$
We use the notation $p\prec q$ for
either ``$p$ interlaces $q$'' or ``$p$ alternates left of $q$''.
The next theorem is a version of the classical Hermite-Biehler theorem.
\begin{theorem}[{\cite[Theorem~4.1]{Branden11}}]\label{Hermite}
Let $F(x)=F^E(x^2)+xF^O(x^2)\in\R[x]$ be standard.
Then $F(x)$ is stable if and only if both $F^E(x)$ and $F^O(x)$ are standard,
have only nonpositive zeros, and $F^O(x)\prec F^E(x)$.
\end{theorem}

The Hermite-Biehler theorem has been widely used to study the distribution of zeros. In this paper, we consider combinatorial aspects of the expression $F(x)=F^E(x^2)+xF^O(x^2)$, which we call the {\it Hermite-Biehler decomposition} of $F(x)$.
In Section~\ref{Section02}, we collect the Hermite-Biehler decompositions of several enumerative polynomials,
including flag descent polynomials for hyperoctahedral group, flag ascent-plateau polynomials for Stirling permutations, up-down run polynomials for symmetric group and
alternating run polynomials for hyperoctahedral group. As applications of Theorem~\ref{mathm01}, we get some new properties of associated polynomials.
\section{Eulerian recurrence systems and Hermite-Biehler decompositions}\label{Section02}
\subsection{The flag descent polynomials}
\hspace*{\parindent}

It follows from~\cite[Theorem 4.3]{Adin2001} that
the flag descent polynomials $C_n(x)$ satisfy the recurrence
\begin{equation}\label{AdinSnx-recu}
C_{n+1}(x)=(2nx^2+x+1)C_n(x)+x(1-x^2)C_n'(x),
\end{equation}
with $C_0(x)=1,C_1(x)=1+x$ and $C_2(x)=1+3x+3x^2+x^3$.
Set $$\alpha_n(x)=2nx^2+x+1,~\beta_n(x)=x(1-x^2).$$ Note that $\beta_n^e(x)=0$.
It follows from~\eqref{dualrelation01} that
$$x(2nx^2+x+1)=xp_n(x^2)+r_n(x^2),~2x^2(1-x^2)=q_n(x^2).$$
Hence $p_n(x)=2nx+1,~q_n(x)=2x(1-x),~r_n(x)=x$.
By Theorem~\ref{mathm01}, we see that
$$u_n(x)=2nx+1+1-x=(2n-1)x+2,~v_n(x)=2x(1-x),~w_n(x)=1.$$
By Theorem~\ref{mathm01} and Theorem~\ref{Hermite}, we get the second main result of this paper.
\begin{theorem}\label{Snx-thm}
For $n\geq1$, we have
\begin{equation}\label{Snx-recu}
\left\{
  \begin{array}{l}
    C_{n+1}^E(x)=(2nx+1)C_n^E(x)+2x(1-x)\frac{d}{dx}{C_n^E}(x)+xC_n^O(x), \\
    C_{n+1}^O(x)=((2n-1)x+2)C_n^O(x)+2x(1-x)\frac{d}{dx}C_n^O(x)+C_n^E(x),
  \end{array}
\right.
\end{equation}
with the initial conditions $C_1^E(x)=C_1^O(x)=1$. Moreover, both $C_n^E(x)$ and $C_n^O(x)$ have only real negative zeros and $C_n^O(x)\prec C_n^E(x)$.
\end{theorem}

Define
$$C^E(x,z)=1+\sum_{n=1}^\infty C_n^E(x)\frac{z^n}{n!},~C^O(x,z)=\sum_{n=1}^\infty C_n^O(x)\frac{z^n}{n!}.$$
By rewriting~\eqref{Snx-recu} in terms of generating functions, we have
\begin{equation}\label{Snx-recu-GF}
\left\{
  \begin{array}{l}
    (1-2xz)\frac{\partial}{\partial z}C^E(x,z)=C^E(x,z)+2x(1-x)\frac{\partial}{\partial x}C^E(x,z)+xC^O(x,z),\\
      (1-2xz)\frac{\partial}{\partial z}C^O(x,z)=(2-x)C^O(x,z)+2x(1-x)\frac{\partial}{\partial x}C^O(x,z)+C^E(x,z).
  \end{array}
\right.
\end{equation}
It is routine to check that
\begin{equation}\label{SnE-GF}
C^E(x,z)=\frac{x-e^{(x-1)z}}{x-e^{2(x-1)z}},~C^O(x,z)=\frac{1-e^{(x-1)z}}{e^{2(x-1)z}-x}.
\end{equation}

It is well known that the exponential generating functions of $A_n(x)$ and $B_n(x)$ are given as follows (see~\cite[Theorem~3.4]{Brenti94}):
\begin{equation*}\label{An-GF}
A(x,z):=\sum_{n=0}^\infty A_n(x)\frac{z^n}{n!}=\frac{x-1}{x-e^{(x-1)z}}=\frac{(1-x)e^{(1-x)z}}{1-xe^{(1-x)z}},
\end{equation*}
\begin{equation*}\label{An-GF}
B(x,z):=\sum_{n=0}^\infty B_n(x)\frac{z^n}{n!}=\frac{(1-x)e^{(1-x)z}}{1-xe^{2(1-x)z}}.
\end{equation*}

We can now present the third main result of this paper.
\begin{theorem}\label{mainthm03}
We have
$$\frac{A(x,2z)}{A(x,z)}=C^E(x,z),~\frac{B(x,z)}{A(x,z)}=1+xC^O(x,z).$$
\end{theorem}
\begin{proof}
Note that
$$A(x,2z)=\frac{x-1}{x-e^{2(x-1)z}}=\frac{x-e^{(x-1)z}}{x-e^{2(x-1)z}}\frac{x-1}{x-e^{(x-1)z}}=C^E(x,z)A(x,z),$$
which yields the first formula. Since
\begin{equation*}
1+xC^O(x,z)=\frac{1-xe^{(1-x)z}}{1-xe^{2(1-x)z}},
\end{equation*}
it follows that
$$B(x,z)=\frac{(1-x)e^{(1-x)z}}{1-xe^{2(1-x)z}}=\frac{1-xe^{(1-x)z}}{1-xe^{2(1-x)z}}\frac{(1-x)e^{(1-x)z}}{1-xe^{(1-x)z}}=\left(1+xC^O(x,z)\right)A(x,z).$$
This completes the proof.
\end{proof}
An immediate consequence of Theorem~\ref{mainthm03} is the following corollary.
\begin{corollary}
For $n\geq 0$, we have
$$2^nA_n(x)=\sum_{k=0}^n\binom{n}{k}A_k(x)C_{n-k}^E(x),$$
$$B_n(x)=A_n(x)+x\sum_{k=0}^{n-1}\binom{n}{k}A_k(x)C_{n-k}^O(x).$$
\end{corollary}

In~\cite{Bagno04}, Bagno and Garber introduced the definition of flag excedance of signed permutations.
Let $\pi\in\mbn$.
The {\it flag excedance number} of $\pi$ is defined by
$$\fexc(\pi):=2\#\{i\in [n]:~\pi(i)>i\}+\#\{i\in [n]:~\pi(i)<0\}.$$
Let $\exc_A(\pi)=\#\{i\in [n]:~\pi(i)>i\}$.
Then $\fexc(\pi)=2\exc_A(\pi)+\negg(\pi)$.
It is well known that
\begin{equation}\label{fdesfexc}
\sum_{\pi\in\mbn}x^{\fdes(\pi)}=\sum_{\pi\in\mbn}x^{\fexc(\pi)},
\end{equation}
which has been extended to colored permutations and affine Weyl groups.
Ordinary and $q$-generalizations of~\eqref{fdesfexc} have been pursued by several authors, see~\cite{Foata11,Mongelli13,Zeng16} for instance.
For example, Mongelli~\cite{Mongelli13} derived some basic properties of the flag excedance polynomials of classical and affine Weyl groups. In particular,
Mongelli~\cite[p.~1221-1222]{Mongelli13} presented a combinatorial proof of~\eqref{fdesfexc}. Recall that $\mdn$ is the Coxeter group of type $D$.
According to~\cite[Corollary~5.1]{Mongelli13},
\begin{equation}\label{Dnfexc}
\sum_{\pi\in\mdn}x^{\fexc(\pi)}=\frac{1}{2}\left((1+x)^nA_n(x)+(1-x)^nA_n(-x)\right).
\end{equation}

By~\eqref{fdesfexc}, we see that $$C_n(x)=\sum_{\pi\in\mdn}x^{\fexc(\pi)}+\sum_{\pi\in\mbn\setminus\mdn}x^{\fexc(\pi)}.$$
Let $C(x,z):=\sum_{n=0}^\infty C_n(x)\frac{z^n}{n!}$. It follows from~\eqref{fdesAnx} that $C(x,z)=A(x,(1+x)z)$.
By using~\eqref{Dnfexc}, we immediately get
\begin{equation*}
\sum_{n=0}^\infty \sum_{\pi\in\mdn}x^{\fexc(\pi)}\frac{z^n}{n!}=\frac{1}{2}\left(C(x,z)+C(-x,z)\right)=C^E(x^2,z).
\end{equation*}
Thus
\begin{equation*}
\sum_{n=0}^\infty \sum_{\pi\in\mbn\setminus \mdn}x^{\fexc(\pi)}\frac{z^n}{n!}=xC^O(x^2,z).
\end{equation*}
So we get the following result, which gives a connection between Coxeter groups $\mbn$ and $\mdn$.
\begin{proposition}\label{CnxBnxdes}
Let $C_n^+=\{\pi\in\mbn:~\pi(1)>0\}$. Then we have
\begin{equation*}
\sum_{\pi\in\mdn}x^{\fexc(\pi)}=\sum_{\pi\in C_n^+}x^{2\des_A(\pi)}.
\end{equation*}
\end{proposition}
A combinatorial proof of Proposition~\ref{CnxBnxdes} would be interesting.
\subsection{Flag ascent-plateau polynomials for Stirling permutations}
\hspace*{\parindent}

Stirling permutations were introduced by Gessel and Stanley~\cite{Gessel78}.
A {\it Stirling permutation} of order $n$ is a permutation of the multiset $\{1,1,2,2,\ldots,n,n\}$ such that
for each $i$, $1\leq i\leq n$, all entries between the two occurrences of $i$ are larger than $i$.
The reader is referred to~\cite{Bona08,Haglund12,Ma17,Ma2001} for the recent study on Stirling permutations.

Denote by $\mqn$ the set of {\it Stirling permutations} of order $n$.
Let $\sigma=\sigma_1\sigma_2\cdots\sigma_{2n}\in\mqn$.
The {\it ascent-plateau number}, {\it left ascent-plateau number} and {\it flag ascent-plateau number} of $\sigma$ are respectively defined by
\begin{align*}
\ap(\pi)&=\#\{i\in[2,2n-1]:~\pi(i-1)<\pi(i)=\pi(i+1)\},\\
\lap(\pi)&=\#\{i\in[2n-1]:~\pi(i-1)<\pi(i)=\pi(i+1),~\pi(0)=0\},\\
\fap(\pi)&=\left\{
               \begin{array}{ll}
                 2\ap(\sigma)+1, & \hbox{if $\sigma_1=\sigma_2$;} \\
                 2\ap(\sigma), & \hbox{otherwise.}
               \end{array}
             \right.
\end{align*}
Clearly, $\fap(\sigma)=\ap(\sigma)+\lap(\sigma)$.
The {\it flag ascent-plateau polynomials} $L_n(x)$ are defined by $$L_n(x)=\sum_{\sigma\in\mqn}x^{\fap(\sigma)}.$$
The polynomials $L_n(x)$ satisfy the recurrence relation
\begin{equation}\label{Fnx-recu}
L_{n+1}(x)=(x+2nx^2)L_n(x)+x(1-x^2)L_n'(x),
\end{equation}
with $L_0(x)=1$ (see~\cite[p.~14]{Ma2001}). The first few $L_n(x)$ are
\begin{align*}
L_1(x)&=x,\\
L_2(x)&=x+x^2+x^3,\\
L_3(x)&=x+3x^2+7x^3+3x^4+x^5.
\end{align*}
Set $\mqn=\mqn^+\cup\mqn^-$, where $\mqn^+=\{\sigma\in\mqn:~\sigma_1<\sigma_2\}$ and $\mqn^-=\{\sigma\in\mqn:~\sigma_1=\sigma_2\}$.
Then
\begin{equation}\label{Lnx-express}
L_n(x)=L_n^E(x^2)+xL_n^O(x^2)=\sum_{\sigma\in\mqn^+}x^{2\ap(\sigma)}+x\sum_{\sigma\in\mqn^-}x^{2\ap(\sigma)}.
\end{equation}

From~\eqref{Fnx-recu}, we see that $\alpha_n(x)=x+2nx^2,~\beta_n(x)=x(1-x^2)$ and $\beta_n^e(x)=0$.
Using~\eqref{dualrelation01}, we obtain $x(x+2nx^2)=xp_n(x^2)+r_n(x^2)$ and $2x^2(1-x^2)=q_n(x^2)$.
Hence $$p_n(x)=2nx,~q_n(x)=2x(1-x),~r_n(x)=x,$$
$$u_n(x)=2nx+1-x=(2n-1)x+1,~v_n(x)=2x(1-x),~w_n(x)=1.$$
Therefore, by Theorem~\ref{mathm01}, we can now present the fouth main result of this paper.
\begin{theorem}
For $n\geq1$, we have
\begin{equation}\label{Tnx-recu}
\left\{
  \begin{array}{l}
    L_{n+1}^E(x)=2nxL_n^E(x)+2x(1-x)\frac{d}{dx}{L_n^E}(x)+xL_n^O(x), \\
    L_{n+1}^O(x)=((2n-1)x+1)L_n^O(x)+2x(1-x)\frac{d}{dx}L_n^O(x)+L_n^E(x),
  \end{array}
\right.
\end{equation}
with the initial conditions $L_1^E(x)=0$ and $L_1^O(x)=1$.
\end{theorem}

The {\it ascent-plateau polynomials} and {\it left ascent-plateau polynomials} are defined by
$$M_n(x)=\sum_{\sigma\in\mqn}x^{\ap(\pi)},~N_n(x)=\sum_{\sigma\in\mqn}x^{\lap(\pi)}.$$
According to~\cite[Theorem 2,~Theorem 3]{Ma15}, we have
\begin{equation*}\label{Mxt}
M(x,t)=\sum_{n\geq 0}M_n(x)\frac{t^n}{n!}=\sqrt{\frac{x-1}{x-e^{2t(x-1)}}},
\end{equation*}
\begin{equation*}\label{Nxt}
N(x,t)=\sum_{n\geq 0}N_n(x)\frac{t^n}{n!}=\sqrt{\frac{1-x}{1-xe^{2t(1-x)}}}.
\end{equation*}
Clearly, $M_n(x)=x^nN_n\left(\frac{1}{x}\right)$. Moreover, according to~\cite[Proposition~1]{Ma17}, we have
\begin{align*}
2^nxA_n(x)&=\sum_{i=0}^n\binom{n}{i}N_i(x)N_{n-i}(x),\\
B_n(x)&=\sum_{i=0}^n\binom{n}{i}N_i(x)M_{n-i}(x).
\end{align*}
It is well known that $A_n(x)$ and $B_n(x)$ are both unimodal (see~\cite{Athanasiadis17} for instance).
Motivated by the above convolution formulas, it is natural to explore the unimodality of $M_n(x)$ and $N_n(x)$.

Let $f(x)=\sum_{i=0}^nf_ix^i$.
We say that $f(x)$ is {\it unimodal} if there exists an index $0\leq k\leq n$ such that $f_0\leq f_1\leq \cdots\leq f_k\geq f_{k+1}\geq\cdots \geq f_n$.
Such an index $k$ is called a {\it mode} of $f(x)$. In the past decades, it is a fundamental problem to determine the location of modes of a sequence
of combinatorial polynomials, see~\cite{Chen09,Wang2005} for instance.
Following~\cite[Definition~2.9]{Schepers13},
we say that $f(x)$ is {\it alternatingly increasing} if
$$f_0\leq f_n\leq f_1\leq f_{n-1}\leq\cdots f_{\lrf{\frac{n+1}{2}}}.$$
Clearly, alternatingly increasing property is a stronger property than unimodality. Very recently,
there has been much work on the alternatingly increasing property of Ehrhart polynomials (see~\cite{Beck2019,Branden18,Solus19}).
In the rest part of this subsection, we shall show that
both $M_n(x)$ and $N_n(x)$ are alternatingly increasing.

If $f(x)$ is symmetric and $\deg f(x)=n$, then it can be expanded as
$$f(x)=\sum_{k=0}^{\lrf{{n}/{2}}}\gamma_kx^k(1+x)^{n-2k},$$ and it is said to be {\it $\gamma$-positive}
if $\gamma_k\geq 0$ for $0\leq k\leq \lrf{\frac{n}{2}}$.
If $f(x)$ is $\gamma$-positive, then $f(x)$ is unomidal and symmetric.
Following~\cite[Definition~15]{Ma2001},
if $$f(x)=(1+x)^\nu\sum_{k=0}^{n}\lambda_kx^k(1+x^2)^{n-k}$$ and $\lambda_k\geq 0$ for all $0\leq k\leq n$,
then we say that $f(x)$ is {\it semi-$\gamma$-positive}, where $\nu=0$ or $\nu=1$.
Therefore, if $f(x)$ is semi-$\gamma$-positive, then
\begin{align*}
f(x)&=(1+x)^\nu\left(\sum_{k=0}^{\lrf{n/2}}\lambda_{2k}x^{2k}(1+x^2)^{n-2k}+\sum_{k=0}^{\lrf{(n-1)/2}}\lambda_{2k+1}x^{2k+1}(1+x^2)^{n-1-2k}\right)\\
&=(1+x)^\nu(f_1(x^2)+xf_2(x^2)),
\end{align*}
and both $f_1(x)$ and $f_2(x)$ are $\gamma$-positive (see~\cite[Proposition~16]{Ma2001}).

We now recall a very recent result on the flag ascent-plateau polynomials.
\begin{proposition}[{\cite[Theorem~19]{Ma2001}}]\label{Lnx-prop}
The polynomial $L_n(x)$ is semi-$\gamma$-positive. More precisely, for $n\geq 1$,
we have $$L_n(x)=\sum_{k=1}^nL_{n,k}x^k(1+x^2)^{n-k},$$
where the numbers $L_{n,k}$ satisfy the recurrence relation
\begin{equation}\label{Lnk-recu}
L_{n+1,k}=kL_{n,k}+L_{n,k-1}+4(n-k+2)L_{n,k-2},
\end{equation}
with the initial conditions $L_{0,0}=1$ and $L_{n,0}=0$ for $n\geq 1$.
\end{proposition}

It follows from~\eqref{Lnx-express} that
\begin{equation}\label{MnxNnx}
M_n(x)=L_n^E(x)+L_n^O(x),~N_n(x)=L_n^E(x)+xL_n^O(x).
\end{equation}
Since $\deg L_n(x)=2n-1$, we have $\deg L_n^E(x)=\deg L_n^O(x)=n-1$.
Note that $L_n^E(0)=0$ and $L_n^O(0)=1$.
By using Proposition~\ref{Lnx-prop}, we get that both $L_n^E(x)$ and $L_n^O(x)$ are $\gamma$-positive for any $n\geq 1$.
More precisely, we have
\begin{equation}\label{LnELnO}
\left\{
  \begin{array}{l}
L_n^E(x)=\sum_{\sigma\in\mqn^+}x^{\ap(\sigma)}=\sum_{k=1}^{\lrf{n/2}} L_{n,2k} x^k(1+x)^{n-2k},, \\
L_n^O(x)=\sum_{\sigma\in\mqn^-}x^{\ap(\sigma)}=\sum_{k=0}^{\lrf{(n-1)/2}} L_{n,2k+1}x^k(1+x)^{n-1-2k}.
  \end{array}
\right.
\end{equation}
%
In conclusion, we now present the fifth main result of this paper.
\begin{theorem}
For any $n\geq 1$, both $M_n(x)$ and $N_n(x)$ are alternatingly increasing, and so they are unimodal with modes in the middle.
\end{theorem}
\begin{proof}
It follows from~\eqref{LnELnO} that both $L_n^E(x)$ and $L_n^O(x)$ are symmetric and unimodal. When $n=2m+1$,
assume that
\begin{align*}
L_{2m+1}^E(x)&=\ell_1x+\ell_2x^2+\cdots+\ell_{m-1}x^{m-1}+\ell_mx^m+\ell_{m}x^{m+1}+\ell_{m-1}x^{m+2}+\cdots \ell_2x^{2m-1}+\ell_1x^{2m},\\
L_{2m+1}^O(x)&=1+\widetilde{\ell}_1x+\widetilde{\ell}_2x^2+\cdots+\widetilde{\ell}_{m-1}x^{m-1}+\widetilde{\ell}_mx^m+\widetilde{\ell}_{m-1}x^{m+1}+\cdots+\widetilde{\ell}_1x^{2m-1}+x^{2m}.
\end{align*}
Then $M_{2m+1}(x)=L_{2m+1}^E(x)+L_{2m+1}^O(x)=\sum_{i=0}^{2m}M_{2m+1,i}x^i$,
where
$$M_{2m+1,i}=\left\{
  \begin{array}{ll}
    1, & \hbox{if $i=0$;} \\
    \ell_i+\widetilde{\ell}_i, & \hbox{if $1\leq i\leq m$;} \\
    \ell_{2m-i+1}+\widetilde{\ell}_{2m-i}, & \hbox{if $m+1\leq i\leq 2m-1$;} \\
    \ell_1+1, & \hbox{if $i=2m$.}
  \end{array}
\right.$$
It is clear that $1\leq \ell_1+1\leq \ell_1+\widetilde{\ell}_1\leq \ell_2+\widetilde{\ell}_1\leq \cdots\leq \ell_m+\widetilde{\ell}_m$, i.e.,
$$M_{2m+1,0}\leq M_{2m+1,2m}\leq M_{2m+1,1}\leq M_{2m+1,2m-1}\leq \cdots\leq M_{2m+1,m}.$$
Note that $N_{2m+1}(x)=L_{2m+1}^E(x)+xL_{2m+1}^O(x)=\sum_{i=0}^{2m+1}N_{2m+1,i}x^i$,
where
$$N_{2m+1,i}=\left\{
  \begin{array}{ll}
    0, & \hbox{if $i=0$;} \\
    \ell_1+1, & \hbox{if $i=1$;} \\
    \ell_i+\widetilde{\ell}_{i-1}, & \hbox{if $2\leq i\leq m$;} \\
    \ell_{2m-i+1}+\widetilde{\ell}_{2m-i+1}, & \hbox{if $m+1\leq i\leq 2m$;} \\
    1, & \hbox{if $i=2m+1$.}
  \end{array}
\right.$$
It is clear that $0<1\leq \ell_1+1\leq \ell_1+\widetilde{\ell}_1\leq \ell_2+\widetilde{\ell}_1\leq \cdots\leq \ell_m+\widetilde{\ell}_m$, i.e.,
$$N_{2m+1,0}\leq N_{2m+1,2m+1}\leq N_{2m+1,1}\leq N_{2m+1,2m}\leq \cdots\leq N_{2m+1,m+1}.$$
Therefore, both $M_{2m+1}(x)$ and $N_{2m+1}(x)$ are alternatingly increasing, the mode of $M_{2m+1}(x)$ is $m$ and that of $N_{2m+1}(x)$ is $m+1$.
In the same way, one can verify that
both $M_{2m}(x)$ and $N_{2m}(x)$ are alternatingly increasing with the mode $m$. This completes the proof.
\end{proof}
\subsection{The up-down run polynomials for symmetric group}
\hspace*{\parindent}

Let $\pi\in\msn$. The {\it up-down runs} of a permutation $\pi\in\msn$ are the alternating runs of $\pi$ endowed with a 0
in the front~(see~\cite{Taylor03,Ma2001}). Let $\udrun(\pi)$ denote the number of up-down runs of $\pi$.
The {\it interior peak} number and {\it left peak} number of $\pi$ are respectively defined by
\begin{align*}
\ipk(\pi)&=\#\{i\in[2,n-1]:~\pi(i-1)<\pi(i)>\pi(i+1)\},\\
\lpk(\pi)&=\#\{i\in[n-1]: \pi(i-1)<\pi(i)>\pi(i+1),~\pi(0)=0\}.
\end{align*}
Define
$$W_n(x)=\sum_{\pi\in\msn}x^{\ipk(\pi)},~\overline{W}_n(x)=\sum_{\pi\in\msn}x^{\lpk(\pi)}.$$
The polynomials $W_n(x)$ and $\overline{W}_n(x)$ satisfy the recurrence rleations
$$W_{n+1}(x)=(nx-x+2)W_n(x)+2x(1-x)W_n'(x),$$
$$\overline{W}_{n+1}(x)=(nx+1)\overline{W}_n(x)+2x(1-x)\overline{W}_n'(x),$$
with $W_1(x)=\overline{W}_1(x)=1$ (see~\cite{Ma121,Petersen06}).
Note that $\deg \overline{W}_n(x)\geq \deg W_n(x)$. Then we
set $$p_n(x)=nx+1,~q_n(x)=2x(1-x),~r_n(x)=0.$$ Note that
$p_n(x)+\frac{1}{2x}q_n(x)=nx-x+2$. Then by using Theorem~\ref{mathm01}, we can define
\begin{align*}
\alpha_n(x)&=p_n(x^2)+\frac{1}{x}r_n(x^2)=nx^2+1,~
\beta_n(x)=\frac{1}{2x}q_n(x^2)=x(1-x^2).
\end{align*}
Therefore, we immediately get the following result, which has been proved in~\cite{Ma121}.
\begin{proposition}
Let $\{R_n(x)\}_{n\geq 1}$ be a sequence of polynomials defined by
\begin{equation}\label{Rnx01}
R_{n+1}(x)=(nx^2+1)R_n(x)+x(1-x^2)R_n'(x),
\end{equation}
with $R_1(x)=1+x$. Then $R_n(x)=\overline{W}_n(x^2)+xW_n(x^2)$.
\end{proposition}

The {\it up-down run polynomials} $T_n(x)$ are defined by $$T_n(x)=\sum_{\pi\in\msn}x^{\udrun(\pi)}.$$
The polynomials $T_n(x)$ satisfy the recurrence relation
\begin{equation}\label{Rnx02}
T_{n+1}(x)=x(nx+1)T_{n}(x)+x\left(1-x^2\right)T_{n}'(x),
\end{equation}
with initial conditions $T_0(x)=1$ and $T_1(x)=x$ (see~\cite{Ma132,Stanley08}).
Comparing~\eqref{Rnx01} with~\eqref{Rnx02}, it is easy to check that $$R_n(x)=\frac{1+x}{x}T_{n}(x).$$

Set $\msn^+=\{\pi\in\msn:~\pi(n-1)>\pi(n)\}$ and $\msn^-=\{\pi\in\msn:~\pi(n-1)<\pi(n)\}$.
We define
$$T_n^E(x)=\sum_{\pi\in\msn^+}x^{\lpk(\pi)},~T_n^O(x)=\sum_{\pi\in\msn^-}x^{\lpk(\pi)}.$$
By definition, we get the following result.
\begin{proposition}\label{PropUnx}
For $n\geq 1$, we have
\begin{equation*}
\left\{
  \begin{array}{l}
T_n(x)=T_n^E(x^2)+xT_n^O(x^2),\\
\overline{W}_n(x)=T_n^E(x)+T_n^O(x).
  \end{array}
\right.
\end{equation*}
\end{proposition}

From~\eqref{Rnx02}, we see that $\alpha_n(x)=nx^2+x,~\beta_n(x)=x(1-x^2)$ and $\beta_n^e(x)=0$.
Using~\eqref{dualrelation01}, we obtain $x(nx^2+x)=xp_n(x^2)+r_n(x^2)$ and $2x^2(1-x^2)=q_n(x^2)$.
Hence $$p_n(x)=nx,~q_n(x)=2x(1-x),~r_n(x)=x,$$
$$u_n(x)=nx+1-x=(n-1)x+1,~v_n(x)=2x(1-x),~w_n(x)=1.$$

By using~\eqref{Rnx02} and~\cite[Theorem~2]{Ma2008}, we see that $T_n(x)$ have only real nonpositive zeros.
Combining Theorem~\ref{mathm01} and Theorem~\ref{Hermite}, we obtain the sixth main result of this paper.
\begin{theorem}
For $n\geq1$, we have
\begin{equation}\label{Tnx-recu}
\left\{
  \begin{array}{l}
    T_{n+1}^E(x)=nxT_n^E(x)+2x(1-x)\frac{d}{dx}{T_n^E}(x)+xT_n^O(x), \\
    T_{n+1}^O(x)=((n-1)x+1)T_n^O(x)+2x(1-x)\frac{d}{dx}T_n^O(x)+T_n^E(x),
  \end{array}
\right.
\end{equation}
with the initial conditions $T_1^E(x)=0$ and $T_1^O(x)=1$.  Moreover, both $T_n^E(x)$ and $T_n^O(x)$ have only real negative zeros and $T_n^O(x)\prec T_n^E(x)$.
\end{theorem}
\subsection{Alternating run polynomials for signed permutations}
\hspace*{\parindent}

Let $\pi\in\mbn$. The peak number and valley number of $\pi$ are respectively defined by
\begin{align*}
\pk(\pi)&=\#\{i\in[n-1]:~\pi(i-1)<\pi(i)>\pi(i+1),~\pi(0)=0\},\\
\val(\pi)&=\#\{i\in[n-1]: \pi(i-1)>\pi(i)<\pi(i+1),~\pi(0)=0\}.
\end{align*}
Recall that $C_n^+=\{\pi\in\mbn:~\pi(1)>0\}$.
We define
\begin{equation}\label{PnxVnx-def}
\begin{split}
U_n(x)&=\sum_{\pi\in C_n^+}x^{\pk(\pi)},~
V_n(x)=\sum_{\pi\in C_n^+}x^{\val(\pi)}.
\end{split}
\end{equation}

According to~\cite[Corollary~7]{Chow14}, the polynomials $U_n(x)$ and $V_n(x)$ satisfy the following system of Eulerian recurrences:
\begin{equation*}
\left\{
  \begin{array}{l}
U_{n+1}(x)=(2nx+1)U_{n}(x)+4x(1-x)U_{n}'(x)+xV_{n}(x),\\
V_{n+1}(x)=(2nx-2x+3)V_{n}(x)+4x(1-x)V_{n}'(x)+U_{n}(x),
  \end{array}
\right.
\end{equation*}
with $U_{0}(x)=1$ and $V_{0}(x)=0$. Note that $\deg U_n(x)\geq \deg V_n(x)$. Thus $(U_n(x),V_n(x))$ is a Eulerian pair.
Put
$$p_n(x)=2nx+1,~q_n(x)=4x(1-x),~r_n(x)=x,$$
$$u_n(x)=2nx-2x+3,~v_n(x)=4x(1-x),~w_n(x)=1.$$
Then $\alpha_n(x)=p_n(x^2)+\frac{1}{x}r_n(x^2)=2nx^2+x+1$ and
$\beta_n(x)=\frac{1}{2x}q_n(x^2)=2x(1-x^2)$.
Then by using Theorem~\ref{mathm01}, we immediately get the following result.
\begin{proposition}
Let $\{H_n(x)\}_{n\geq 0}$ be a sequence of polynomials defined by
\begin{equation}\label{Rnx01}
H_{n+1}(x)=(2nx^2+x+1)H_n(x)+2x(1-x^2)H_n'(x),
\end{equation}
with $H_0(x)=1$. Then $H_n(x)=U_n(x^2)+xV_n(x^2)$.
\end{proposition}

Let $\altrun(\pi)=\pk(\pi)+\val(\pi)$ be the number of alternating runs of $\pi$.
The alternating run polynomials for signed permutations are given as follows:
$$\widetilde{H}_n(x)=\sum_{\pi\in C_n^+}x^{\altrun(\pi)}.$$
Zhao~\cite{Zhao11} showed that
$\widetilde{H}_{n+1}(x)=(2nx^2+3x-1)\widetilde{H}_{n}(x)+2x\left(1-x^2\right)\widetilde{H}_{n}'(x)$ for $n\geqslant 1$,
with $\widetilde{H}_1(x)=x$. It is routine to verify that $$H_n(x)=\frac{1+x}{x}\widetilde{H}_n(x)$$ for $n\geq 1$, which has been proved in~\cite[Theorem~8]{Chow14}.
%
%
\section{Concluding remark}
In this paper, we consider the combinatorial aspects of the Hermite-Biehler decompositions of several enumerative polynomials.
Let $\{f_n(x)\}_{n\geq 0}$ be a sequence of polynomials with nonnegative coefficients. Suppose that
\begin{equation}\label{recufnx}
f_{n+1}(x)=\left(a_1n+a_2+(b_1n+b_2)x+(c_1n+c_2)x^2\right)f_n(x)+dx(1-x^2)f_n'(x).
\end{equation}
where $a_1,a_2,b_1,b_2,c_1,c_2,d\in\mathbb{R}$. Then
$$\alpha_n(x)=a_1n+a_2+(b_1n+b_2)x+(c_1n+c_2)x^2,~\beta_n(x)=dx(1-x^2).$$
It follows from~\eqref{dualrelation01} that
$$p_n(x)=a_1n+a_2+(c_1n+c_2)x,~q_n(x)=2dx(1-x),~r_n(x)=(b_1n+b_2)x.$$
By using Theorem~\ref{mathm01}, we obtain
$$u_n(x)=a_1n+a_2+d+(c_1n+c_2-d)x,~v_n(x)=2dx(1-x),~w_n(x)=b_1n+b_2,$$
and then we can derive the recurrence system of the polynomials $f^E(x)$ and $f^O(x)$.

Besides the polynomials discussed in Section~\ref{Section02}, many other enumerative polynomials also satisfy the recurrence~\eqref{recufnx}, see~\cite{Barbero14,Zhu,Zhuang17} for instance.
We end this paper by giving an example. Following~\cite[Definition~1]{Aval13}, a tree-like tableau is a Ferrers diagram where each cell
contains either 0 or 1 point with some constraints. The symmetric tableaux are tree-like tableaux which are invariant with respect to reflection
through the main diagonal of their diagram.
Let $b(n,k)$ be the number of symmetric tableaux of size $2n+1$ with $k$ diagonal cells, and let $b_n(x)=\sum_{k=1}^{n+1}b(n,k)x^k$.
It follows from~\cite[Proposition~18]{Aval13} that $$b_{n+1}(x)=(n+1)x(1+x)b_n(x)+x(1-x^2)b_n'(x),$$
with the initial condition $b_0(x)=x$. By using the recurrence system of the polynomials $b_n^E(x)$ and $b_n^O(x)$,
one can easily derive that $b_n^E(1)=b_n^O(1)=2^{n-1}n!$ for $n\geq 1$. We leave the details to the reader.
Thus it may be interesting to further explore properties of $b_n^E(x)$ and $b_n^O(x)$.
\section*{Acknowledgements.}
This work is supported by NSFC (12071063,11571235) and
NSC (108-2115-M-017-005-MY2,~107-2115-M-001-009-MY3).

\end{document}